# Black-Box $k$-to-$1$-PCA Reductions: Theory and Applications

### v2

@@ Abstract
The $k$-principal component analysis ($k$-PCA) problem is a fundamental algorithmic primitive that is widely-used in data analysis and dimensionality reduction applications. In statistical settings, the goal of $k$-PCA is to identify a top eigenspace of the covariance matrix of a distribution, which we only have implicit access to via samples. Motivated by these implicit settings, we analyze black-box deflation methods as a framework for designing $k$-PCA algorithms, where we model access to the unknown target matrix via a black-box $1$-PCA oracle which returns an approximate top eigenvector, under two popular notions of approximation. Despite being arguably the most natural reduction-based approach to $k$-PCA algorithm design, such black-box methods, which recursively call a $1$-PCA oracle $k$ times, were previously poorly-understood.
Our main contribution is significantly sharper bounds on the approximation parameter degradation of deflation methods for $k$-PCA. For a quadratic form notion of approximation we term ePCA (energy PCA), we show deflation methods suffer no parameter loss. For an alternative well-studied approximation notion we term cPCA (correlation PCA), we tightly characterize the parameter regimes where deflation methods are feasible. Moreover, we show that in all feasible regimes, $k$-cPCA deflation algorithms suffer no asymptotic parameter loss for any constant $k$. We apply our framework to obtain state-of-the-art $k$-PCA algorithms robust to dataset contamination, improving prior work both in sample complexity and approximation quality.
@@ 1 Introduction
Principal component analysis (PCA) stands as a ubiquitous technique in the areas of statistical analysis and dimensionality reduction (see e.g. [Pea01] and the popular reference [JC16]), offering a powerful and general-purpose means to extract information from complex datasets. In one of its most common use cases, applied to a distribution $\mathcal{D}$, the $k$-PCA problem (Definition 4) seeks to identify $k$ orthogonal principal components (PCs) which capture the most variation in $\mathcal{D}$. Succinctly put, the goal is to output the top-$k$ eigenvectors of the covariance $\bm{\Sigma}$ of $\mathcal{D}$, [^1] The standard definition of an exact $k$-PCA allows for arbitrary tiebreaking if the dimensionality of the top-$k$ eigenspace is $>k$, see Definition 4 for a formal statement. or an approximation thereof, where the key challenge is that we only have implicit sample access to $\bm{\Sigma}$.
This statistical $k$-PCA problem is extremely well-studied, and due to its fundamental nature, many recent works have explored the development of $k$-PCA algorithms under various constraints. These variations include performing PCA in small space [${}^{+}$16], under adversarial corruptions [JLT20, DKPP23], when the distribution is heavy-tailed [Min15, WM17], or when the sampling process admits correlations [NSW23, KS23]. In some problem settings, additional constraints might be imposed on the data distribution (e.g. sparsity [ZHT06]) or the output of the PCA algorithm (e.g. differential privacy [LKJO22], fairness [LCYY23]). In all these scenarios, asking for an exact $k$-PCA is impossible due to our implicit access to the covariance matrix $\bm{\Sigma}$ in question, so the goal is to output an approximate $k$-PCA, for appropriate notions of approximation.
@@
One of the most commonly-proposed solutions to alleviate the burden of $k$-PCA algorithm design, assuming existence of a corresponding $1$-PCA algorithm, is a reduction-based approach, known as a deflation method [Mac08, AZL16] (see Algorithm 1). In this framework, the extraction of a single approximate PC is treated as a subroutine, and the deflation method repeatedly projects out the results of calls to this subroutine from the dataset, $k$ times in total. It is straightforward to see that if the subroutine returned an exact top eigenvector each time, this strategy would indeed succeed in outputting an exact $k$-PCA of the target matrix. As mentioned previously, under implicit access the key challenge is to quantify the approximation degradation of this recursive procedure.
Indeed, despite their being perhaps the most natural reduction-based $k$-PCA strategy, there has been surprisingly-limited work rigorously analyzing the performance of deflation methods. For example, standard analyses require strong gap assumptions on the target matrix (e.g. each of the top-$k$ eigenvalues being separated) [LZ15], and this condition has even been quoted in a variety of works as being necessary [LZ15, MM15, Sha16b, LKO22]. The challenge of obtaining provable guarantees for deflation methods (or small-block variants thereof) was an open problem stated in [MM15]. Our primary contribution is a direct analysis of the approximation parameter degradation of deflation methods, with no gap assumptions on the target matrix. This refutes the aforementioned conventional wisdom that deflation methods fail without strong spectral assumptions.
Our work is motivated by a closely-related result of [AZL16], in the white-box setting where the target matrix is explicitly available. [^2] For example, in this white-box setting, one can query matrix-vector products with the target matrix, which is not a realistic assumption if the target matrix is the covariance of a distribution we have sample access to. However, the analysis of [AZL16] required the single-PC calls to be highly-accurate, losing large polynomial factors (or worse) in the approximation quality parameters. Due to this overhead, the [AZL16] result does not apply to the implicit black-box setting that we study, which is typically of interest in statistical applications, where one would require an enormous sample complexity to obtain such accurate single-PC calls.
@@ 1.1 Our results
We model black-box access to a target matrix $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$ through the template in Algorithm 1, a black-box $k$-to-$1$-PCA reduction. Specifically, the algorithm only interacts with $\mathbf{M}$ through an oracle $\mathcal{O}_{1\textup{PCA}}$, which we assume returns an approximate top eigenvector of $\mathbf{P}\mathbf{M}\mathbf{P}$ for a specified projection matrix $\mathbf{P}$. This also explains why Algorithm 1 is often called a deflation method, as it repeatedly “deflates” directions from consideration as specified by recursively calling $\mathcal{O}_{1\textup{PCA}}$.
1 Input: $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, $k\in[d]$, $\mathcal{O}_{1\textup{PCA}}$, an algorithm which takes as input $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$ and a $d\times d$ orthogonal projection matrix $\mathbf{P}$ and returns a unit vector in $\mathbb{R}^{d}$ in $\operatorname{span}(\mathbf{P})$; $\mathbf{P}_{0}\leftarrow\mathbf{I}_{d}$ for $i\in[k]$ do
2       $u_{i}\leftarrow\mathcal{O}_{1\textup{PCA}}(\mathbf{M},\mathbf{P}_{i-1})$ $\mathbf{P}_{i}\leftarrow\mathbf{P}_{i-1}-u_{i}u_{i}^{\top}$
@@
If $\mathcal{O}_{1\textup{PCA}}$ is assumed to be an exact $1$-PCA algorithm, then Algorithm 1 is an exact $k$-PCA algorithm. In Sections 3 and 4, we analyze the behavior and approximation quality of Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ only has approximate $1$-PCA guarantees. We consider two types of approximation which are well-studied in the literature, which we call energy PCA (or ePCA, see Definition 1) and correlation PCA (or cPCA, see Definition 2). Correspondingly, our main results concern Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ is assumed to be an approximate $1$-ePCA oracle or $1$-cPCA oracle, notions formalized in Definitions 5 and 6, and we want Algorithm 1 to respectively return a $k$-ePCA or a $k$-cPCA of $\mathbf{M}$.
@@
$\left\langle\mathbf{U}\mathbf{U}^{\top},\mathbf{M}\right\rangle\geq(1-\epsilon)\left\lVert\mathbf{M}\right\rVert_{k}.$
Our guarantees for Algorithm 1 as an ePCA algorithm are stated in the following result.
@@
The other notion of approximation we consider is correlation $k$-PCA, a popular definition in the literature [Sha16a, AZL17] which defines the performance of an approximate PCA algorithm in terms of the correlation with the small eigenspace of $\mathbf{M}$, allowing for a gap in the definition of “small.” Let $\mathbf{V}^{<\lambda}(\mathbf{M})$ denote the orthonormal matrix spanning the eigenspace of $\mathbf{M}$ corresponding to eigenvalues $<\lambda$ [^3] For notation used throughout the paper, see Section 2. We define cPCA as follows, parameterized by a gap $\Gamma$ and a correlation $\Delta$.
@@
Definition 2 requires that $\mathbf{U}$ be almost entirely-contained in an eigenspace of $\mathbf{M}$ corresponding to eigenvalues which are, at worst, barely outside the top-$k$ space. We remark that if $\mathbf{U}$ is an exact $k$-PCA of $\mathbf{M}$ (Definition 4), it is also a 00-$k$-ePCA and a $(0,0)$-$k$-cPCA. Before we state our main result on cPCA, we point out that not all values of $\left(\delta,\gamma\right)$ pairs [^4] We sometimes use lowercase Greek letters to denote parameters for the $1$-cPCA oracle, to differentiate from capital Greek letters used to parameterize overall $k$-PCA guarantees. are feasible inputs for Algorithm 1 to serve as a black-box reduction, unlike in the case of ePCA. Formally, we characterize two types of parameter regimes for black-box cPCA using the following definition.
Definition 3 (Black-box cPCA regimes).
$\delta:=\frac{\Delta}{f(k)},\;\gamma:=\frac{\Gamma}{f(k)},$ (1)
Proposition 1 (Invalid black-box cPCA regimes).
We establish Proposition 1 in Section 4.2 by giving $3$-dimensional examples where Algorithm 1 immediately fails to return a $(\Delta,\Gamma)$-$2$-cPCA if $\Gamma\lesssim\kappa_{2}(\mathbf{M})\sqrt{\Delta}$ asymptotically, assuming $\mathcal{O}_{1\textup{PCA}}$ is a $(\delta,\gamma)$-$1$-cPCA oracle, for any $\delta,\gamma$ smaller than $\Delta,\Gamma$ by a constant.
This begs the natural question: are the bounds on $\alpha$ and $\beta$ in Proposition 1 tight? We answer affirmatively with our main cPCA result in Theorem 2, proven in Section 4, showing that if $\Delta\cdot\kappa_{k}(\mathbf{M})^{2}\leq\Gamma^{2}$, then Algorithm 1 indeed serves as a black-box cPCA reduction. Proposition 1 and Theorem 2 hence completely classify black-box cPCA regimes, up to the dependence on $k$.
@@
In other words, for constant $k$, if $\Gamma^{2}\gtrsim\Delta\cdot\kappa_{k}(\mathbf{M})^{2}$, then Algorithm 1 is lossless asymptotically without any gap assumptions on $\mathbf{M}$, and otherwise it provably does not serve as a generic cPCA reduction even for $k=2$. We mention that our choice to parameterize the regimes $g$ in Definition 3 by $\kappa_{k}(\mathbf{M})$ is motivated by a result of [Lia23] (see Eq. (1.3) in that paper), which proves that some sample complexity overhead in $\kappa_{k}(\mathbf{M})$ is necessary in general. On the other hand, the requirement that $\Gamma\gtrsim\sqrt{\Delta}$ appears to be an artifact of the specific deflation strategy in Algorithm 1.

### v3

@@ Abstract
The $k$-principal component analysis ($k$-PCA) problem is a fundamental algorithmic primitive that is widely-used in data analysis and dimensionality reduction applications. In statistical settings, the goal of $k$-PCA is to identify a top eigenspace of the covariance matrix of a distribution, which we only have black-box access to via samples. Motivated by these settings, we analyze black-box deflation methods as a framework for designing $k$-PCA algorithms, where we model access to the unknown target matrix via a black-box $1$-PCA oracle which returns an approximate top eigenvector, under two popular notions of approximation. Despite being arguably the most natural reduction-based approach to $k$-PCA algorithm design, such black-box methods, which recursively call a $1$-PCA oracle $k$ times, were previously poorly-understood.
Our main contribution is significantly sharper bounds on the approximation parameter degradation of deflation methods for $k$-PCA. For a quadratic form notion of approximation we term ePCA (energy PCA), we show deflation methods suffer no parameter loss. For an alternative well-studied approximation notion we term cPCA (correlation PCA), we tightly characterize the parameter regimes where deflation methods are feasible. Moreover, we show that in all feasible regimes, $k$-cPCA deflation algorithms suffer no asymptotic parameter loss for any constant $k$. We apply our framework to obtain state-of-the-art $k$-PCA algorithms robust to dataset contamination, improving prior work in sample complexity by a $\textup{poly}(k)$ factor.Accepted for presentation at the Conference on Learning Theory (COLT) 2024
@@ 1 Introduction
Principal component analysis (PCA) stands as a ubiquitous technique in the areas of statistical analysis and dimensionality reduction (see e.g. [Pea01] and the popular reference [JC16]), offering a powerful and general-purpose means to extract information from complex datasets. In one of its most common use cases, applied to a distribution $\mathcal{D}$, the $k$-PCA problem (Definition 1) seeks to identify $k$ orthogonal principal components (PCs) which capture the most variation in $\mathcal{D}$. Succinctly put, the goal is to output the top-$k$ eigenvectors of the covariance $\bm{\Sigma}$ of $\mathcal{D}$, [^1] The standard definition of an exact $k$-PCA allows for arbitrary tiebreaking if the dimensionality of the top-$k$ eigenspace is $>k$, see Definition 1 for a formal statement. or an approximation thereof, where the key challenge is that we only have black-box sample access to $\bm{\Sigma}$.
This statistical $k$-PCA problem is extremely well-studied, and due to its fundamental nature, many recent works have explored the development of $k$-PCA algorithms under various constraints. These variations include performing PCA in small space [JJK+16], under adversarial corruptions [JLT20, DKPP23], when the distribution is heavy-tailed [Min15, WM17], or when the sampling process admits correlations [NSW23, KS23]. In some problem settings, additional constraints might be imposed on the data distribution (e.g. sparsity [ZHT06]) or the output of the PCA algorithm (e.g. differential privacy [LKJO22], fairness [LCYY23]). In all these scenarios, asking for an exact $k$-PCA is impossible due to our black-box access to the covariance matrix $\bm{\Sigma}$ in question, so the goal is to output an approximate $k$-PCA, for appropriate notions of approximation.
@@
One of the most commonly-proposed solutions to alleviate the burden of $k$-PCA algorithm design, assuming existence of a corresponding $1$-PCA algorithm, is a reduction-based approach, known as a deflation method [Mac08, AZL16] (see Algorithm 1). In this framework, the extraction of a single approximate PC is treated as a subroutine, and the deflation method repeatedly projects out the results of calls to this subroutine from the dataset, $k$ times in total. It is straightforward to see that if the subroutine returned an exact top eigenvector each time, this strategy would indeed succeed in outputting an exact $k$-PCA of the target matrix. As mentioned previously, under black-box access, the key challenge is to quantify the approximation degradation of this recursive procedure.
Indeed, despite their being perhaps the most natural reduction-based $k$-PCA strategy, there has been surprisingly-limited work rigorously analyzing the performance of deflation methods. For example, standard analyses require strong gap assumptions on the target matrix (e.g. each of the top-$k$ eigenvalues being separated) [LZ15], and this condition has even been quoted in a variety of works as being necessary [LZ15, MM15, Sha16, LKO22]. The challenge of obtaining provable guarantees for deflation methods (or small-block variants thereof) was an open problem stated in [MM15]. Our primary contribution is a direct analysis of the approximation parameter degradation of deflation methods, with no gap assumptions on the target matrix. This refutes the aforementioned conventional wisdom that deflation methods fail without strong spectral assumptions.
Our work is motivated by a closely-related result of [AZL16], which primarily focused on applications in the white-box setting where the target matrix is explicitly available. [^2] For example, in this white-box setting, one can query matrix-vector products with the target matrix, which is not a realistic assumption if the target matrix is the covariance of a distribution we have sample access to. However, the [AZL16] analysis loses large polynomial factors in the approximation quality parameters, under their notion of approximation (Definition 4), which we will shortly compare to our results. Our focus in this work, in comparison, is to give tighter characterizations of when it is possible to obtain lossless or nearly-lossless reductions in the black-box PCA setting, typically of interest in statistical applications.
@@ 1.1 Our results
We model black-box access to a target matrix $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$ through the template in Algorithm 1, a black-box $k$-to-$1$-PCA reduction. [^3] For notation used throughout the paper, see Section 2; $\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$ is the set of positive semidefinite $d\times d$ matrices. Specifically, the algorithm only interacts with $\mathbf{M}$ through an oracle $\mathcal{O}_{1\textup{PCA}}$, which we assume returns an approximate top eigenvector of $\mathbf{P}\mathbf{M}\mathbf{P}$ for a specified projection matrix $\mathbf{P}$. This also explains why Algorithm 1 is often called a deflation method, as it repeatedly “deflates” directions from consideration as specified by recursively calling $\mathcal{O}_{1\textup{PCA}}$.
1 Input: $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, $k\in[d]$, $\mathcal{O}_{1\textup{PCA}}$, an algorithm which takes as input $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$ and a $d\times d$ orthogonal projection matrix $\mathbf{P}$ and returns a unit vector in $\mathbb{R}^{d}$ in $\operatorname{span}(\mathbf{P})$; $\mathbf{P}_{0}\leftarrow\mathbf{I}_{d}$
2 for $i\in[k]$ do
3       $u_{i}\leftarrow\mathcal{O}_{1\textup{PCA}}(\mathbf{M},\mathbf{P}_{i-1})$
4       $\mathbf{P}_{i}\leftarrow\mathbf{P}_{i-1}-u_{i}u_{i}^{\top}$
5
@@ 4       $\mathbf{P}_{i}\leftarrow\mathbf{P}_{i-1}-u_{i}u_{i}^{\top}$
To motivate our approximation-tolerant analysis, first consider the performance of Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ returns an exact $1$-PCA of $\mathbf{P}\mathbf{M}\mathbf{P}$ on inputs $(\mathbf{M},\mathbf{P})$, using the following definition.
Definition 1 (Exact PCA).
For $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, the exact $k$-principal component analysis ($k$-PCA) problem asks to return orthonormal $\mathbf{V}\in\mathbb{R}^{d\times k}$ such that $\left\langle\mathbf{V}\mathbf{V}^{\top},\mathbf{M}\right\rangle$ is maximal.
If $\mathcal{O}_{1\textup{PCA}}$ is assumed to be an exact $1$-PCA algorithm, then it is straightforward to show that Algorithm 1 is an exact $k$-PCA algorithm (indeed, this follows as a special case of our Theorem 1).
In Sections 3 and 4, our focus is analyzing the behavior of Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ only has approximate $1$-PCA guarantees. We consider two types of approximation which are well-studied in the literature, that we call energy PCA (or ePCA, see Definition 2) and correlation PCA (or cPCA, see Definition 4). We chose these definitions of approximate PCA because of their prevalence as error metrics in the PCA literature; to our knowledge, essentially all works on PCA in the statistical setting (i.e., under sample access from a distribution) give guarantees under one of these metrics, including the aforementioned works [Min15, JJK+16, WM17, JLT20, LKJO22, KS23, DKPP23, LCYY23]. Correspondingly, our main results concern Algorithm 1 when $\mathcal{O}_{1\textup{PCA}}$ is assumed to be an approximate $1$-ePCA oracle or $1$-cPCA oracle, notions formalized in Definitions 3 and 5, and we want Algorithm 1 to respectively return a $k$-ePCA or a $k$-cPCA of $\mathbf{M}$.
@@
$\left\langle\mathbf{U}\mathbf{U}^{\top},\mathbf{M}\right\rangle\geq(1-\epsilon)\left\lVert\mathbf{M}\right\rVert_{k},\text{ where }\left\lVert\mathbf{M}\right\rVert_{k}:=\max_{\textup{orthonormal }\mathbf{V}\in\mathbb{R}^{d\times k}}\left\langle\mathbf{V}\mathbf{V}^{\top},\mathbf{M}\right\rangle.$
Our guarantees for Algorithm 1 as a black-box ePCA reduction are stated in the following result.
Definition 3 (ePCA oracle).
@@
The other notion of approximation we consider is correlation $k$-PCA, a popular definition in the literature [Sha16, AZL17] which defines the performance of an approximate PCA algorithm in terms of the correlation with the small eigenspace of $\mathbf{M}$, allowing for a gap in the definition of “small.” Let $\mathbf{V}^{<\lambda}(\mathbf{M})$ denote the orthonormal matrix spanning the eigenspace of $\mathbf{M}$ corresponding to eigenvalues $<\lambda$. We define cPCA as follows, parameterized by a gap $\Gamma$ and a correlation $\Delta$.
@@
Definition 4 requires that $\mathbf{U}$ be almost entirely-contained in an eigenspace of $\mathbf{M}$ corresponding to eigenvalues which are, at worst, barely outside the top-$k$ space. For example, in the gapped setting when $\lambda_{2}(\mathbf{M})<(1-\Gamma)\lambda_{1}(\mathbf{M})$ and $k=1$ (so $\mathbf{V}^{<(1-\Gamma)\lambda_{k}(\mathbf{M})}$ is simply the subspace orthogonal to the top PC of $\mathbf{M}$), Definition 4 implies recovery of the top PC in the $\sin^{2}$ error metric. In general, cPCA guarantees can be viewed as more geometrically explicit than ePCA counterparts (as they bound correlation with specific eigenspaces), and hence may be preferable when they are available. [^4] However, in some black-box settings e.g. robust PCA (Section 5.2), certificates on specific eigenspaces are harder to achieve, and only weaker certificates such as quadratic forms have found algorithmic use (giving ePCA guarantees).
The only analysis of the black-box deflation method for $k$-PCA (Algorithm 1) we are aware of is due to [AZL16], who gave a black-box $k$-to-$1$-cPCA reduction under the following definition.
Definition 5 (cPCA oracle).
We say $\mathcal{O}_{\textup{cPCA}}$ is a $(\delta,\gamma)$-approximate [^5] We sometimes use lowercase Greek letters to denote parameters for the $1$-cPCA oracle, to differentiate from capital Greek letters used to parameterize overall $k$-PCA guarantees. $1$-cPCA oracle (or, $(\delta,\gamma)$-$1$-cPCA oracle) if, on inputs $\mathbf{M}\in\mathbb{S}_{\succeq\mathbf{0}}^{d\times d}$, and $\mathbf{P}\in\mathbb{R}^{d\times d}$, where $\mathbf{P}$ is required to be an orthogonal projection matrix, $\mathcal{O}_{\textup{cPCA}}$ returns $u\in\mathbb{R}^{d}$, a $(\delta,\gamma)$-$1$-cPCA of $\mathbf{P}\mathbf{M}\mathbf{P}$, satisfying $u\in\operatorname{span}(\mathbf{P})$.
Under Definitions 45, the main result Theorem 4.1(a) of [AZL16] can be rephrased as follows: Algorithm 1 returns a $(\Delta,\Gamma)$-$k$-cPCA if $\mathcal{O}_{1\textup{PCA}}$ is assumed to be a $(\delta,\gamma)$-$1$-cPCA oracle, for
$\gamma=\frac{\Gamma}{2},\;\delta=\Theta\left(\frac{\Delta^{2}\Gamma^{2}}{k^{4}\kappa_{k}(\mathbf{M})^{2}}\right),\text{ where }\kappa_{k}(\mathbf{M}):=\frac{\lambda_{1}(\mathbf{M})}{\lambda_{k}(\mathbf{M})}.$ (1)
In the white-box setting, $(\delta,\gamma)$-$1$-cPCA runtimes typically scale polynomially in $\frac{1}{\gamma}$ and logarithmically in $\frac{1}{\delta}$ (e.g., Theorem 1, [MM15]), so [AZL16] focused on maintaining a low overhead in $\gamma$. However, known sample complexity lower bounds in the black-box statistical setting (e.g., Theorem 6, [AZL17]) show $\gtrsim\frac{1}{\gamma^{2}\delta}$ samples are required to obtain a $(\delta,\gamma)$-$1$-cPCA, even under strong tail bounds such as sub-Gaussianity. Hence, the [AZL16] reduction in the black-box setting uses
$\approx\frac{k^{4}\kappa_{k}(\mathbf{M})^{2}}{\Delta^{2}\Gamma^{4}},$ (2)
samples for obtaining a $(\Delta,\Gamma)$-$k$-cPCA via Algorithm 1, times problem-dependent factors such as the dimension, which is a $\textup{poly}(k,\Delta,\Gamma)$ factor worse than the lower bound. Our goal is to understand whether this parameter loss is inherent for black-box cPCA reductions. [^6] Note that Theorem 1 shows no such loss is necessary, if our approximation metric is instead ePCA. As a starting point, in Section 2, we prove the following simple results transferring guarantees between Definitions 2 and 4.
Lemma 1.
Lemma 2.
By combining Lemmas 1 and 2 with our black-box ePCA reduction in Theorem 1, it is straightforward to show Algorithm 1 returns a $(\Delta,\Gamma)$-$k$-cPCA if $\mathcal{O}_{1\textup{PCA}}$ is a $(\delta,\gamma)$-$1$-cPCA oracle, for
$\delta=\gamma=\Theta\left(\frac{\Delta\Gamma}{k\kappa_{k}(\mathbf{M})}\right).$ (3)
Our transfer lemmas and Theorem 1 therefore already yield a sample complexity scaling as
$\approx\frac{k^{3}\kappa_{k}(\mathbf{M})^{3}}{\Delta^{3}\Gamma^{3}}$ (4)
in standard statistical settings, which is incomparable to (2) in general but already improves upon it in, e.g., the dependences on $k$ and $\Gamma$. However, both (2) and (4) yield polynomial overheads in various approximation factors. We hence ask the natural question: can we characterize when Algorithm 1 serves as an asymptotically lossless cPCA reduction, even when $k$ is a constant?
In Section 4, we give our main result analyzing the use of Algorithm 1 as a black-box $k$-to-$1$-cPCA reduction. We show that if $\Delta\cdot\kappa_{k}(\mathbf{M})^{2}\leq\Gamma^{2}$ for a target matrix $\mathbf{M}$, then Algorithm 1 indeed is an asymptotically lossless reduction for $k=\Theta(1)$. More precisely, we prove the following result.
@@
That is, when $\Delta\cdot\kappa_{k}(\mathbf{M})^{2}\leq\Gamma^{2}$, it suffices to take $\mathcal{O}_{1\textup{PCA}}$ to be a $(\delta,\gamma)$-$1$-cPCA oracle, for $\delta=\Theta_{k}(\Delta)$, $\gamma=\Theta_{k}(\Gamma)$. We complement Theorem 1 with an impossibility result (see Proposition 1), showing that if the condition $\Delta\cdot\kappa_{k}(\mathbf{M})^{2}\leq\Gamma^{2}$ is violated, then Algorithm 1 fails to have this lossless reduction property, even when $k=2$ and $d=3$. In conclusion, Proposition 1 and Theorem 2 completely classify regimes where Algorithm 1 acts as a lossless cPCA reduction, up to the dependence on $k$.